I thought biology was difficult, geez. I don't see any potential solutions for the 1st question. the x's cancel each other out. wait x = -2 by subtracting 2 from both sides gives x = -2?
1.sq root /x + 2 - x = 0 any potential solutions? Cancel out the x’s
This is undefined?
2.
C computes the cost in millions of $ of implementing a city infastructure project when x percent of the citizens participate.
C(x) = 1.5x / 100-x
Using this model, approximately what is the cost if 40% of the citizens participate?
Using this model, estimate the percentage of participation that can be excepted if $5 million is spent on this infastructure project? Set up an equation and solve algebraically. Round to the nearest whole percent.
3.
Suppose that N=squarerootx +2 models the number of cases of an infection, in millions, of a disease x years from now.
a) Approximately how many cases of the infection will there be 16 years from now?
b) In approximately how many years will there be 7 million cases?

Answer by stanbon(75887)

1.sq root /x + 2 - x = 0 any potential solutions?
Is it sqrt(x) or is it sqrt(x+2) ?
====================
2.
C computes the cost in millions of $ of implementing a city infastructure project when x percent of the citizens participate.
C(x) = 1.5x / 100-x
Using this model, approximately what is the cost if 40% of the citizens participate?
C(40) = 1.5*40/(100-40) = 60/60 = 1 million dollars
--------------------------
Using this model, estimate the percentage of participation that can be excepted if $5 million is spent on this infastructure project? Set up an equation and solve algebraically. Round to the nearest whole percent.
5 = 1.5x/(100-x)
500-5x = 1.5x
500 = 6.5x
x = 76.9%
-------------------3.
Suppose that N=sqrt(x) +2 models the number of cases of an infection, in millions, of a disease x years from now.
a) Approximately how many cases of the infection will there be 16 years from now?
N = sqrt(16)+2 = 4+2 = 6 million
-------------------------
b) In approximately how many years will there be 7 million cases?
7 = sqrt(x)+2
5 = sqrt(x)
x = 25 yrs
=========================
Cheers,
Stan H.